# ArguMentor: Augmenting User Experiences with Counter-Perspectives

### v2

@@ Results
Figure 4 shows the number of claims increase by 3, 4, and 5 times when the users use the system for neutral, right-leaning and left-leaning articles, respectively. These differences were significant (p=0.001, 0.0003, 0.0001). The overall word count also increased by over 200% for the claims.
@@
Similar to claims, the number of counter-arguments users were able to generate increased significantly, ranging from a 0.6x increase (neutral) to a 6x increase (left-leaning) (p=0.03, 0.002, 0.0001 for neutral, right- and left-leaning, respectively). Figure 4 shows the number of counter-arguments with and without the system. The overall word count for the counter-arguments again increased by over 200% over the baseline. Detailed diagrams are in the Supplemental Materials.
@@
Figure 5 shows how claims per minute increase for all types of articles when participants use ArguMentor. On the other hand, counter-arguments per minute goes down for neutral articles, and goes up for left- and right-leaning articles. Participants spend more time and generate fewer counter-arguments when reading neutral articles with ArguMentor compared to reading without it.
@@
Figure 6 shows how long participants spend on which feature in the system. This data is acquired using Mouseflow logs. Over half the time was spent on claims and counter-arguments, while the Q&A and DebateMe features were least used with less than 10% of overall time spent. Interestingly, these results align with the preliminary study’s survey results in Table 1, where participants used main claims and counter-arguments the most.
@@
Change in views:
Table 2 shows how user views changed for right-leaning, left-leaning and neutral articles before and after the system. As other studies have previously shown, political views are more entrenched, and harder to change, which is reflected in this study as well. The change in views for neutral articles was significant (p$<$0.05), whereas left and right articles were not.
@@
Political Side Article Topic p-value
R Large corporations should pay a much higher tax than the rest of us. 0.58
R Minimum wage is a good idea 0.75
L Sometimes, violence is the only way to protest. 0.21
L Identity politics is an important idea for social movements to propagate their views 0.12
N Social media does more good than harm 0.03*
N Remote work is a net positive. 0.02*
Table 2: Subjective Opinions Results: Change in Views. * = difference is significant.
@@ Discussion
In this paper, we propose ArguMentor, a system designed to help users gain a more nuanced, balanced understanding of news opinion pieces. We note that users often encounter one perspective when reading news articles, and could benefit from having access to counter-arguments. The larger goal of our system is not necessarily to change user opinions, but to give users quick, easy access to multiple viewpoints while reading an opinion piece, helping them form a more informed, robust, and nuanced perspective.
@@
Addressing RQ1, we can see from the outcomes results that participants identified a significantly higher number of claims and counter-arguments with the help of the system. Without the system, most participants could only identify one main claim within the article. However, the system systematically highlights all claims throughout the article and provides them with specific counters, which can help users holistically understand the article. Participants were instructed not to return to the system or article when filling the survey, but were still able to write 4x as many points after using ArguMentor, showing that the system of highlighting also helps with memory. This is consistent with other HCI systems which use highlighting to boost memory (e.g., (Fok et al. 2022)).
@@
In case of RQ2, we can see that the additional time participants spent on the system often led to a higher number of claims, and counter-arguments. The only exception was common knowledge topics like remote work and social media, where people already were able to come up counter-arguments without the system. However, users are able to generate longer counters with the help of the system (as indicated by a higher average word count), indicating higher confidence and potentially higher quality of counters. Beyond opinion pieces, news articles, especially about economics, science, politics, etc., can be complex, jargon-heavy, and unintentionally biased, indicating places where our system can provide additional value.
For RQ3, we found that ArguMentor led to significant changes in participants’ views on topics, but only for neutral articles. It may be that neutral articles provided more viewpoints and opportunities for exploring counter-arguments and alternative perspectives, or that the topics of these articles (e.g., remote work) were less polarizing. Similar to other works in the field, we speculate that participants didn’t change views for other articles because political views are more ingrained, and harder to change.
@@
When the topic/article was particularly complicated, users used the highlight feature (and benefited from it) a lot more. The right-aligned economics article on taxation, which was naturally harder and jargon-heavier than others, showed an increased usage of highlight features, etc- and users later reported finding those features useful as well. Similarly, even though DebateMe was used less overall, users who were particularly passionate about a news topic used it (and found it useful) a lot more for that article. Users also reported finding the thumbs up and down features useful, as they were able to get several different arguments quite easily.
The results from this system can be used to present news in a different way to users. Journalists working for news companies could take inspiration from ArguMentor to make their straight news as unbiased as possible, whereas op-ed writers could use it to identify and address likely opposing viewpoitns. On social media platforms, AI and LLMs can also be used to automatically generate counter-arguments (like warnings) under some posts to educate and engage users. Humans benefit from exposure to multiple perspectives to be able to make better independent decisions, and ArguMentor’s passive and active support features offer an important step towards that goal.
@@
Additionally, all participants in the study were homogeneous in terms of high formal education and enthusiasm for news. Consequently, their willingness to read opinion pieces, consider multiple perspectives, and engage with system features may differ from a broader cross section of the news-reading public. However, our preliminary study indicated that even this participant pool could find it difficult to carefully read and focus on op-eds. Future work should extend this evaluation to a broader demographic pool.

### v3

@@ Results
Figure 6 shows the number of claims increase by at least two times when the users engage with the system. These differences were significant (p=0.01, 0.001, 0.005). The overall word count also increased by over 200% for the claims. Along with the topic, we also averaged each participant’s improvement using the system over the baseline. A percent increase of 170% is observed overall, which shows that our system helped each individual as well.
@@
Similar to claims, the number of counter-arguments users were able to generate increased significantly, ranging from a 0.6x increase (economics) to a 2.5x increase (social) (p=0.01, 0.004, 0.02) for economics, politics and social, respectively). Figure 6 shows the number of counter-arguments with and without the system. The overall word count for the counter-arguments again increased by over 150% over the baseline. Similar to claims, counter-arguments saw an individual percent increase of 150% across all topics.
@@
Critical Thinking Skills
Figure 6 shows how various critical thinking skills changed with the baseline. Biases, questions and solutions were judged, and all saw an increase with the system, even though the system doesn’t directly do any of this for the user. For the ACT style questions, participants answered, on average, 60% of the questions correctly without the system, and 78% of the questions correctly with the system, showing another indirect benefit of the system.
Figure 6: Performance: Average Critical Thinking Skills over all three domains
Quality of Arguments
The computational quality assessment (Llama-3 7B) preferred the counter-arguments generated by participants after they used ArguMentor 83% of the time. For the human experts, 35/48 instances showed unanimous agreement and hence were kept. Out of this, experts preferred counter-arguments generated by participants after they used the system 71% of the time. This shows that both human experts and AI-as-judge agreed that arguments were higher quality after the participants experienced the system.
@@
Figure 7 shows how claims per minute increase for all types of articles when participants use ArguMentor. Counter-arguments per minute are similar for the baseline and the system.
@@
Figure 8 shows how long participants spend on which feature in the system. This data is acquired using Mouseflow logs. Over half the time was spent on claims and counter-arguments, while the Q&A and DebateMe features were least used with less than 10% of overall time spent. These results align with the preliminary study’s survey results in Table 1, where participants used main claims and counter-arguments the most.
@@ Discussion
In this paper, we propose ArguMentor, a system designed to help users gain a more nuanced, balanced understanding of news opinion pieces. We note that users often encounter one perspective when reading news articles, and could benefit from having access to counter-arguments. The larger goal of our system is to give users quick, easy access to multiple viewpoints while reading an opinion piece, helping them form a more informed, robust, and nuanced perspective.
@@
Addressing RQ1, we can see from the outcomes results that participants identified a significantly higher number of claims and counter-arguments with the help of the system. Without the system, most participants could only identify one main claim within the article. However, the system systematically highlights all claims throughout the article and provides them with specific counters, which can help users holistically understand the article. Participants were instructed not to return to the system or article when filling the survey, but were still able to write up to 3 times as many points after using ArguMentor, showing that the system of highlighting also helps with memory. This is consistent with other HCI systems which use highlighting to boost memory (e.g., (Fok et al. 2022)). This finding was noticed by participants as well; e.g., P4 mentioned how “highlighting was the best source of the system. I didn’t have to even read the entire essay, I got the gist of it quickly.”
We can also see that critical thinking skills rose across the bar for a variety of topics. Participants reported using creative methods like asking more questions, finding out all definitions, etc to identify biases, and pose solutions. They also reported using the counter-arguments to come up with a more nuanced solution to any problems in the article. In the exit interview, multiple participants talked about how reading the counter-arguments made them think of more innovative solutions to the presented problem. P11 talked about his process: ”To detect biases, I mainly used the counter-arguments, then asked questions to the system about how the counter fixes issues in the claims”. P20 talked about how reading counters made him think of his experiences and how both the claim and counter could be improved. An interesting thing we noticed is that most often users didn’t repeat what the AI gave them as ”rebuttals” when writing counter-arguments in the survey, rather adding their own details and thought process to it. This shows that a system like ArguMentor gets the user thinking about a potential counter. Most of our users echoed this, while P16 said ”sometimes I thought the counter-argument was a bit weak, maybe it lacked recent nuance about the elections or something, but it got me thinking about a potential counters that apply to that too”
@@
In case of RQ2, we can see that the additional time participants spent on the system often led to a higher number of claims, and a similar number of counter-arguments. With counter-arguments the time it takes users to read the counter-argument presented by the system potentially adds to the delay, making it roughly equal to the time spent thinking using a baseline. However, users are able to generate longer counters with the help of the system (as indicated by a higher average word count), indicating higher confidence. Moreover, the quality of the counter-argument is higher, as indicated by both human experts and AI-as-judge. Beyond opinion pieces, news articles, especially about economics, science, politics, etc., can be complex, jargon-heavy, and unintentionally biased, indicating places where our system can provide additional value.
For RQ3, we found that ArguMentor was easy to use, helpful, and participants were motivated to use it. However, the recommendation for daily use is lesser because of the time it takes to not only read the article, but also read all the additional information the system provides.
@@
When the topic or article was particularly complicated, participants used ArguMentor’s highlight feature (and benefited from it) more. The economics article on stock buybacks which was naturally harder and jargon-heavier than others, showed an increased usage of highlight features, and participants later reported finding those features useful as well. Similarly, even though DebateMe was used less overall, participants who were particularly passionate about a news topic used it (and found it useful) more for that article. Participants also reported finding the thumbs up and down features useful, as they were able to get several different arguments quite easily.
Impacts on Critical Thinking and Reading Skills
An interesting question this raises, and that future systems should also take into account is whether such a system promotes or erodes critical thinking. On one hand, the results show that the facets of critical thinking (asking questions, identifying biases, writing counter-arguments) are all better with the system. On the other hand, there is not enough evidence to prove whether these are better because there is some inherent change in the ArguMentor user. This forces us to think about the role such a system plays in today’s society. It fills in a few gaps for the users to fill in the rest. Argumentation and reasoning are skills that we develop over our lifetime, and such a system can help with critical reading on a regular basis to then encourage critical thinking even without the system. There is a risk, similar to the risk ChatGPT poses to our education system, where users get accustomed to system generated answers and stop thinking beyond that. To mitigate this risk, this system only gives users tools to think critically. We help them with a few potential rebuttals to claims they find in the passage, so that they can decide what they believe, what biases are present on both sides, and what solutions can come up with a compromise from both sides. We don’t directly provide solutions or biases. Our participants echoed this experience in their exit interview.
Moreover, for those less educated, this tool could provide value in developing critical reading skills. Our participants that had a high school education or less talked about how this system was different from their usual reading experience and how it felt like it aided them: P1 said ”[ArguMentor] didn’t shame me, didn’t make fun of me for not knowing what, [for example] a stock buyback was — that article was a tough read, and I got through it with all the definitions, highlights, I would love something like this everyday.” Even if users completely ignored any counter-arguments or other features, they can simply read definitions of words in the passages and understand it better, something they probably wouldn’t do otherwise. This tool, when used correctly, can enhance literacy on these topics, and reduce misinformation and bias.
@@
Even if poor arguments are generated, which aren’t persuasive to users, at the least it offers users a chance to think of counter-arguments. It encourages the user to confront the possibility that the counter-argument provided is a poor one and that something else could potentially be better.
Conclusion
In this work, we introduce ArguMentor, a novel human-AI collaboration system designed to enhance critical thinking and combat echo chambers in news consumption. ArguMentor highlights claims, identifies counter-arguments, and generates context-based summaries for the user. With active features like a question-answering bot and a debating agent, ArguMentor encourages users to engage deeply with the text. Our evaluation, conducted on a diverse group of participants, shows that users not only generate more claims and higher-quality counter-arguments but also raise more questions, identify biases, and propose solutions more frequently than when reading articles without the system. ArguMentor highlights a creative way in which AI can serve as a tool that boosts rather than erodes critical thinking, learning, and engagement with argumentation.